# On the kernels of the pro-$p$ outer Galois representations associated to once-punctured CM elliptic curves

### v2

@@ 3.2. Two-variable filtration on pro-$p$ mapping class group
We define two subgroups $\tilde{\Gamma}$ and $\Gamma^{{\dagger}}$ of $\mathrm{Aut}(\Pi)$ as
$\tilde{\Gamma}\coloneqq\left\{f\in\mathrm{Aut}(\Pi)\,\middle|\,\begin{aligned}&\text{$\bar{f}$ preserves $\langle\bar{x}\rangle$ and $\langle\bar{y}\rangle$ respectively, and}\\ &\text{$f$ preserves the conjugacy class of $\langle z\rangle$}\end{aligned}\right\}\text{and}$
@@
where $\bar{f}\in\mathrm{Aut}(\Pi^{\mathrm{ab}})$ is the automorphism of $\Pi$ induced by $f$ in. Moreover, we set $\Gamma\coloneqq\tilde{\Gamma}/\mathrm{Inn}(\Pi)$, which is a subgroup of $\mathrm{Out}(\Pi)$.
One can easily observe that the subgroup $\Pi(\boldsymbol{m})$ defined in the last subsection is preserved under the action of $\tilde{\Gamma}$. By the definition of $\tilde{\Gamma}$, there exist two homomorphisms [^3] This notation is ambiguous since we already use the symbols $\chi_{1}$ and $\chi_{2}$ to indicate the characters $G_{K}\to\mathbb{Z}_{p}^{\times}$. However, if we identify $\Pi$ with the fundamental group $\Pi_{1,1}$ via $x=x_{1}$ and $y=x_{2}$, then the image of the pro-$p$ outer Galois representation $\rho_{X,p}$ is contained in $\Gamma$, and these notations become compatible.
$\chi_{1}\colon\tilde{\Gamma}\to\mathrm{Aut}(\langle\bar{x}\rangle)=\mathbb{Z}_{p}^{\times}\quad\text{and}\quad\chi_{2}\colon\tilde{\Gamma}\to\mathrm{Aut}(\langle\bar{y}\rangle)=\mathbb{Z}_{p}^{\times}$
which are characterized by the equalities $\bar{f}(\bar{x})=\bar{x}^{\chi_{1}(f)}$ and $\bar{f}(\bar{y})=\bar{y}^{\chi_{2}(f)}$ for every $f\in\tilde{\Gamma}$, respectively. Note that these characters factor through the quotient $\Gamma$. There is also a homomorphism $\Gamma^{{\dagger}}\to\mathrm{Aut}(\langle z\rangle)=\mathbb{Z}_{p}^{\times}$ by the definition of $\Gamma^{{\dagger}}$, which is found to coincide with the product $\chi_{1}\chi_{2}$ since the commutator map $\Pi/\Pi(2)\times\Pi/\Pi(2)\to\Pi(2)/\Pi(3)$ is bilinear.
Remark 3.3.
If we identify $\Pi$ with the pro-$p$ fundamental group $\Pi_{1,1}$ of the once-punctured CM elliptic curve $X=E-O$ via $x=x_{1}$ and $y=x_{2}$, then the group $\tilde{\Gamma}$ (resp. $\Gamma^{{\dagger}}$, $\Gamma$) is a subgroup of $\tilde{\Gamma}_{1,1}$ (resp. $\Gamma_{1,1}^{{\dagger}}$, $\Gamma_{1,1}$) introduced in Section 2.1. Moreover, through this identification, the action of the absolute Galois group $G_{K}$ preserves $\langle\bar{x}\rangle$ and $\langle\bar{y}\rangle$, since they are isomorphic to the $\mathfrak{p}$-adic and the $\bar{\mathfrak{p}}$-adic Tate module of $E$, respectively. In particular, the image of the pro-$p$ outer Galois representation $\rho_{X,p}:G_{K}\to\Gamma_{1,1}$ is contained in $\Gamma$.
@@
$F^{m}\tilde{\Gamma}\coloneqq\ker\left(\tilde{\Gamma}\to\mathrm{Aut}(\Pi/\Pi(m+1))\right)$
@@
$F^{\boldsymbol{m}}\tilde{\Gamma}\coloneqq\ker\left(\tilde{\Gamma}\to\prod_{\boldsymbol{k}\in\{(0,1),(1,0)\}}\mathrm{Aut}\left(\Pi(\boldsymbol{k})/\Pi(\boldsymbol{m}+\boldsymbol{k})\right)\right),$
$F^{\boldsymbol{m}}\Gamma^{{\dagger}}\coloneqq F^{\boldsymbol{m}}\Gamma\cap\Gamma^{{\dagger}},\quad\text{and}\quad F^{\boldsymbol{m}}\Gamma\coloneqq\mathrm{Im}(F^{\boldsymbol{m}}\tilde{\Gamma}\to\Gamma).$
Note that, since $[\Pi(\boldsymbol{m}),\Pi(\boldsymbol{k})]\subset\Pi(\boldsymbol{m}+\boldsymbol{k})$ for $\boldsymbol{k}\in\{(1,0),(0,1)\}$, the inner automorphism group $\mathrm{Inn}_{\Pi(\boldsymbol{m})}(\Pi)$ of $\Pi$ induced by elements of $\Pi(\boldsymbol{m})$ is contained in $F^{\boldsymbol{m}}\tilde{\Gamma}$.
@@
Here, we use $[f(y)^{-n}y^{n},f(x)]\in[\Pi(\boldsymbol{m}+(0,1)),\Pi(1,0)]\subset\Pi(\boldsymbol{m}+\boldsymbol{1})\subset\Pi(\boldsymbol{m}+(1,0))$ to establish the first congruence. A similar computation shows that $f$ also acts trivially on $\Pi(0,1)/\Pi(\boldsymbol{m}+(0,1))$. ∎
@@
Let $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$. Then the group $F^{\boldsymbol{m}}\tilde{\Gamma}$ acts trivially on the quotient $\Pi(\boldsymbol{n})/\Pi(\boldsymbol{n}+\boldsymbol{m})$ for every $\boldsymbol{n}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$.
@@
We prove the assertion by induction on $\lvert\boldsymbol{n}\rvert$. First, if $\lvert\boldsymbol{n}\rvert=1$, the assertion follows by the definition of $F^{\boldsymbol{m}}\tilde{\Gamma}$. In general, since $\Pi(\boldsymbol{n})$ is generated by subgroups $[\Pi(\boldsymbol{n}^{\prime}),\Pi(\boldsymbol{n}^{\prime\prime})]$ such that $\boldsymbol{n}^{\prime}+\boldsymbol{n}^{\prime\prime}=\boldsymbol{n}$ and $0<|\boldsymbol{n}^{\prime}|,|\boldsymbol{n}^{\prime\prime}|<|\boldsymbol{n}|$, the assertion follows by the fact that the commutator map
@@
For $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, we define $\mathrm{gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}$ and $\mathrm{gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}$ by
$\mathrm{gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}\coloneqq F^{\boldsymbol{m}}\tilde{\Gamma}/F^{\boldsymbol{m}+(1,0)}\tilde{\Gamma}\quad\text{and}\quad\mathrm{gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}\coloneqq F^{\boldsymbol{m}}\tilde{\Gamma}/F^{\boldsymbol{m}+(0,1)}\tilde{\Gamma}.$
Similarly, we define $\mathrm{gr}^{\boldsymbol{m}}_{1}\Gamma$ and $\mathrm{gr}^{\boldsymbol{m}}_{2}\Gamma$ by
$\mathrm{gr}^{\boldsymbol{m}}_{1}\Gamma\coloneqq F^{\boldsymbol{m}}\Gamma/F^{\boldsymbol{m}+(1,0)}\Gamma\quad\text{and}\quad\mathrm{gr}^{\boldsymbol{m}}_{2}\Gamma\coloneqq F^{\boldsymbol{m}}\Gamma/F^{\boldsymbol{m}+(0,1)}\Gamma.$
@@
$\mathrm{gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}\to\mathrm{gr}^{\boldsymbol{m}}_{1}\Gamma\quad\text{and}\quad\mathrm{gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}\to\mathrm{gr}^{\boldsymbol{m}}_{2}\Gamma.$
@@
$i_{\boldsymbol{m},1}\colon\mathrm{gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}\to\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{1}\Pi\oplus\mathrm{gr}^{\boldsymbol{m}+(0,1)}_{1}\Pi\colon f\mapsto(f(x)x^{-1},f(y)y^{-1})\quad\text{and}$
$i_{\boldsymbol{m},2}\colon\mathrm{gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}\to\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{2}\Pi\oplus\mathrm{gr}^{\boldsymbol{m}+(0,1)}_{2}\Pi\colon f\mapsto(f(x)x^{-1},f(y)y^{-1}).$
@@
$\Gamma^{{\dagger}}_{1}\coloneqq\tilde{\Gamma}_{1}\cap\Gamma^{{\dagger}},\quad\Gamma_{1}\coloneqq\mathrm{Im}(\tilde{\Gamma}_{1}\to\Gamma),$
$\Gamma^{{\dagger}}_{2}\coloneqq\tilde{\Gamma}_{2}\cap\Gamma^{{\dagger}},\quad\Gamma_{2}\coloneqq\mathrm{Im}(\tilde{\Gamma}_{2}\to\Gamma).$
Then the action of $\tilde{\Gamma}_{1}/F^{(1,0)}\tilde{\Gamma}$ (resp. $\tilde{\Gamma}_{2}/F^{(0,1)}\tilde{\Gamma}$) on $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ (resp. $\mathrm{gr}^{\boldsymbol{m}}_{2}\Pi$) commutes with the action of $\Pi/\Pi(1,0)$ (resp. $\Pi/\Pi(0,1)$).
In the following, we study the action of the quotient $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ (resp. $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}$) on the graded quotient $\mathrm{gr}^{\boldsymbol{m}}_{1}$ (resp. $\mathrm{gr}^{\boldsymbol{m}}_{2}$) of the groups $\Pi$, $\tilde{\Gamma}$ and $\Gamma$.
@@
For every $\gamma\in\Gamma^{{\dagger}}_{1}$, we have $\gamma(x)x^{-\chi_{1}(\gamma)}\in\Pi(2,0)$. Similarly, the element $\gamma(y)y^{-\chi_{2}(\gamma)}$ is contained in $\Pi(0,2)$ for every $\gamma\in\Gamma^{{\dagger}}_{2}$.
@@
We only prove the first assertion, since the proof of the second one is similar. Since $\gamma\in\Gamma^{{\dagger}}_{1}$ acts $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-linearly on $\Pi(1,0)/\Pi(2,0)$, which is a free $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-module generated by $x$, the action of $\gamma$ is given by a scalar multiplication by an element of $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]^{\times}$. However, since we have $z=[y,x]=(y-1)x\neq 0$ in $\Pi(1,0)/\Pi(2,0)$ and $\gamma(z)=z^{\chi_{1}(\gamma)}$, such a scalar coincides with $\chi_{1}(\gamma)$. Hence we have $\gamma(x)\equiv x^{\chi_{1}(\gamma)}\bmod\Pi(2,0)$ as desired. ∎
@@
Let $\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$. Then the action of the group $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ on the graded quotient $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ is given by the character $\chi_{1}^{m_{1}}$. Similarly, the action of $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}$ on $\mathrm{gr}^{\boldsymbol{m}}_{2}\Pi$ is given by $\chi_{2}^{m_{2}}$.
@@
Fix an element $\gamma\in\Gamma^{{\dagger}}_{1}$, and we prove by induction on $|\boldsymbol{m}|$ that the action of $\gamma$ on $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ is given by multiplication by $\chi_{1}(\gamma)^{m_{1}}$. If $|\boldsymbol{m}|=1$, the assertion follows from (the proof of) Lemma 3.7. In general, the graded quotient $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ is generated (as a $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-module) by the image of commutator maps
$[\;\cdot\;,\;\cdot\;]\colon\mathrm{gr}^{\boldsymbol{m}^{\prime}}_{1}\Pi\times\mathrm{gr}^{\boldsymbol{m}^{\prime\prime}}_{1}\Pi\to\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi,$
where $\boldsymbol{m}^{\prime}=(m_{1}^{\prime},m_{2}^{\prime})$ and $\boldsymbol{m}^{\prime\prime}=(m_{1}^{\prime\prime},m_{2}^{\prime\prime})$ satisfy $\boldsymbol{m}^{\prime}+\boldsymbol{m}^{\prime\prime}=\boldsymbol{m}$. Since this pairing is bilinear, for every $(\tau^{\prime},\tau^{\prime\prime})\in\mathrm{gr}^{\boldsymbol{m}^{\prime}}_{1}\Pi\times\mathrm{gr}^{\boldsymbol{m}^{\prime\prime}}_{1}\Pi$, we have
@@
$i_{\boldsymbol{m},1}\colon\mathrm{gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}\to\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{1}\Pi(\chi_{1}^{-1})\oplus\mathrm{gr}^{\boldsymbol{m}+(0,1)}_{1}\Pi$
is compatible with the action of $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$. Here,x $\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{1}\Pi(\chi_{1}^{-1})$ denotes the $\chi_{1}^{-1}$-twist of $\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{1}\Pi$. Similarly, the homomorphism
$i_{\boldsymbol{m},2}\colon\mathrm{gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}\to\mathrm{gr}^{\boldsymbol{m}+(1,0)}_{2}\Pi\oplus\mathrm{gr}^{\boldsymbol{m}+(0,1)}_{2}\Pi(\chi_{2}^{-1})$
is compatible with the action of $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}$. In particular, the action of $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ (resp. $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}$) on $\mathrm{gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}$ (resp. $\mathrm{gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}$) is given by $\chi_{1}^{m_{1}}$ (resp. $\chi_{2}^{m_{2}}$).
@@
Before proving the assertion, we note that the group $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ acts on the graded quotients $\mathrm{gr}^{\boldsymbol{m}}_{1}\Pi$ by the character $\chi_{1}^{m_{1}}$, by Lemma 3.8. Take an arbitrary element $\gamma$ of $\Gamma^{{\dagger}}_{1}$. We prove the first assertion by computing the image of $\gamma\cdot f$ under the map $i_{\boldsymbol{m},1}$ for an arbitrary $f\in F^{\boldsymbol{m}}\tilde{\Gamma}$.
@@
Since there exists $\alpha\coloneqq y^{-1}\gamma^{-1}(y)$ is contained in $\Pi(2)=\Pi(\boldsymbol{1})$, we have

### v3

@@ 3.2. Two-variable filtration on pro-$p$ mapping class group
We define two subgroups $\tilde{\Gamma}$ and $\Gamma^{{\dagger}}$ of $\operatorname{Aut}(\Pi)$ as
[…]
@@
where $\bar{f}\in\operatorname{Aut}(\Pi^{\mathrm{ab}})$ denotes the automorphism of $\Pi^{\mathrm{ab}}$ induced by $f$. We set $\Gamma\coloneqq\tilde{\Gamma}/\operatorname{Inn}(\Pi)$, which is a subgroup of $\operatorname{Out}(\Pi)$.
One can easily observe that the subgroup $\Pi(\boldsymbol{m})$ defined in the last subsection is preserved under the action of $\tilde{\Gamma}$. Moreover, by the definition of $\tilde{\Gamma}$, there exist two characters [^2] This notation is ambiguous since we already use the symbols $\chi_{1}$ and $\chi_{2}$ to indicate the characters $G_{K}\to\mathbb{Z}_{p}^{\times}$. However, if we identify $\Pi$ with the fundamental group $\Pi_{1,1}$ via $x=x_{1}$ and $y=x_{2}$, then the image of the pro-$p$ outer Galois representation $\rho_{X,p}$ is contained in $\Gamma$, and these notations become compatible.
$\chi_{1}\colon\tilde{\Gamma}\to\operatorname{Aut}(\langle\bar{x}\rangle)=\mathbb{Z}_{p}^{\times}\quad\text{and}\quad\chi_{2}\colon\tilde{\Gamma}\to\operatorname{Aut}(\langle\bar{y}\rangle)=\mathbb{Z}_{p}^{\times}$
which are characterized by the equalities $\bar{f}(\bar{x})=\bar{x}^{\chi_{1}(f)}$ and $\bar{f}(\bar{y})=\bar{y}^{\chi_{2}(f)}$. These characters factor through the quotient $\Gamma$. There is also a character $\Gamma^{{\dagger}}\to\operatorname{Aut}(\langle z\rangle)=\mathbb{Z}_{p}^{\times}$ by the definition of $\Gamma^{{\dagger}}$, but this coincides with the product $\chi_{1}\chi_{2}$, since the commutator map $\Pi/\Pi(2)\times\Pi/\Pi(2)\to\Pi(2)/\Pi(3)$ is bilinear.
Remark 3.6.1.
If we identify $\Pi$ with the pro-$p$ fundamental group $\Pi_{1,1}$ of the once-punctured CM elliptic curve $X$ via $x=x_{1}$ and $y=x_{2}$, then $\tilde{\Gamma}$ (resp. $\Gamma^{{\dagger}}$, $\Gamma$) is a subgroup of $\tilde{\Gamma}_{1,1}$ (resp. $\Gamma_{1,1}^{{\dagger}}$, $\Gamma_{1,1}$) introduced in Section 2.1. Moreover, through this identification, the $G_{K}$-action preserves $\langle\bar{x}\rangle$ and $\langle\bar{y}\rangle$, since they are isomorphic to the $\mathfrak{p}$-adic and the $\bar{\mathfrak{p}}$-adic Tate module of $E$, respectively. In particular, the image of the pro-$p$ outer Galois representation $\rho_{X,p}:G_{K}\to\Gamma_{1,1}$ is contained in $\Gamma$.
@@
$F^{m}\tilde{\Gamma}\coloneqq\operatorname{Ker}\left(\tilde{\Gamma}\to\operatorname{Aut}(\Pi/\Pi(m+1))\right)$
@@
$F^{\boldsymbol{m}}\tilde{\Gamma}\coloneqq\operatorname{Ker}\left(\tilde{\Gamma}\to\prod_{\boldsymbol{k}\in\{(0,1),(1,0)\}}\operatorname{Aut}\left(\Pi(\boldsymbol{k})/\Pi(\boldsymbol{m}+\boldsymbol{k})\right)\right),$
$F^{\boldsymbol{m}}\Gamma^{{\dagger}}\coloneqq F^{\boldsymbol{m}}\Gamma\cap\Gamma^{{\dagger}},\quad\text{and}\quad F^{\boldsymbol{m}}\Gamma\coloneqq\operatorname{Im}(F^{\boldsymbol{m}}\tilde{\Gamma}\to\Gamma).$
Note that, since $\Pi(\boldsymbol{m}),\Pi(\boldsymbol{k})]\subset\Pi(\boldsymbol{m}+\boldsymbol{k})$ for $\boldsymbol{k}\in\{(1,0),(0,1)\}$, the inner automorphism group $\operatorname{Inn}_{\Pi(\boldsymbol{m})}(\Pi)$ of $\Pi$ induced by elements of $\Pi(\boldsymbol{m})$ is contained in $F^{\boldsymbol{m}}\tilde{\Gamma}$.
@@
Here, we use $f(y)^{-n}y^{n},f(x)]\in[\Pi(\boldsymbol{m}+(0,1)),\Pi(1,0)]\subset\Pi(\boldsymbol{m}+\boldsymbol{1})\subset\Pi(\boldsymbol{m}+(1,0))$ to establish the first congruence. A similar computation shows that $f$ also acts trivially on $\Pi(0,1)/\Pi(\boldsymbol{m}+(0,1))$. ∎
@@
Let $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$. Then $F^{\boldsymbol{m}}\tilde{\Gamma}$ acts trivially on $\Pi(\boldsymbol{n})/\Pi(\boldsymbol{n}+\boldsymbol{m})$ for every $\boldsymbol{n}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$.
@@
We prove the assertion by induction on $\lvert\boldsymbol{n}\rvert$. First, if $\lvert\boldsymbol{n}\rvert=1$, the assertion follows by the definition of $F^{\boldsymbol{m}}\tilde{\Gamma}$. In general, since $\Pi(\boldsymbol{n})$ is generated by subgroups $\Pi(\boldsymbol{n}^{\prime}),\Pi(\boldsymbol{n}^{\prime\prime})]$ such that $\boldsymbol{n}^{\prime}+\boldsymbol{n}^{\prime\prime}=\boldsymbol{n}$ and $0<|\boldsymbol{n}^{\prime}|,|\boldsymbol{n}^{\prime\prime}|<|\boldsymbol{n}|$, the assertion follows by the fact that the commutator map
@@
For $\boldsymbol{m}\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$, we define $\operatorname{Gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}$ and $\operatorname{Gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}$ by
$\operatorname{Gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}\coloneqq F^{\boldsymbol{m}}\tilde{\Gamma}/F^{\boldsymbol{m}+(1,0)}\tilde{\Gamma}\quad\text{and}\quad\operatorname{Gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}\coloneqq F^{\boldsymbol{m}}\tilde{\Gamma}/F^{\boldsymbol{m}+(0,1)}\tilde{\Gamma}$.
Similarly, we define $\operatorname{Gr}^{\boldsymbol{m}}_{1}\Gamma$ and $\operatorname{Gr}^{\boldsymbol{m}}_{2}\Gamma$ by
$\operatorname{Gr}^{\boldsymbol{m}}_{1}\Gamma\coloneqq F^{\boldsymbol{m}}\Gamma/F^{\boldsymbol{m}+(1,0)}\Gamma\quad\text{and}\quad\operatorname{Gr}^{\boldsymbol{m}}_{2}\Gamma\coloneqq F^{\boldsymbol{m}}\Gamma/F^{\boldsymbol{m}+(0,1)}\Gamma.$
@@
$\operatorname{Gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}\to\operatorname{Gr}^{\boldsymbol{m}}_{1}\Gamma\quad\text{and}\quad\operatorname{Gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}\to\operatorname{Gr}^{\boldsymbol{m}}_{2}\Gamma.$
@@
$i_{\boldsymbol{m},1}\colon\operatorname{Gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}\to\operatorname{Gr}^{\boldsymbol{m}+(1,0)}_{1}\Pi\oplus\operatorname{Gr}^{\boldsymbol{m}+(0,1)}_{1}\Pi\colon f\mapsto(f(x)x^{-1},f(y)y^{-1})\quad\text{and}$
$i_{\boldsymbol{m},2}\colon\operatorname{Gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}\to\operatorname{Gr}^{\boldsymbol{m}+(1,0)}_{2}\Pi\oplus\operatorname{Gr}^{\boldsymbol{m}+(0,1)}_{2}\Pi\colon f\mapsto(f(x)x^{-1},f(y)y^{-1}).$
@@
$\Gamma^{{\dagger}}_{1}\coloneqq\tilde{\Gamma}_{1}\cap\Gamma^{{\dagger}},\quad\Gamma_{1}\coloneqq\operatorname{Im}(\tilde{\Gamma}_{1}\to\Gamma),$
$\Gamma^{{\dagger}}_{2}\coloneqq\tilde{\Gamma}_{2}\cap\Gamma^{{\dagger}},\quad\Gamma_{2}\coloneqq\operatorname{Im}(\tilde{\Gamma}_{2}\to\Gamma).$
Then the action of $\tilde{\Gamma}_{1}/F^{(1,0)}\tilde{\Gamma}$ (resp. $\tilde{\Gamma}_{2}/F^{(0,1)}\tilde{\Gamma}$) on $\operatorname{Gr}^{\boldsymbol{m}}_{1}\Pi$ (resp. $\operatorname{Gr}^{\boldsymbol{m}}_{2}\Pi$) commutes with the action of $\Pi/\Pi(1,0)$ (resp. $\Pi/\Pi(0,1)$). In the following, we study the action of $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ (resp. $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}$) on $\operatorname{Gr}^{\boldsymbol{m}}_{1}$ (resp. $\operatorname{Gr}^{\boldsymbol{m}}_{2}$) of various groups.
@@
The element $\gamma(x)x^{-\chi_{1}(\gamma)}$ is contained in $\Pi(2,0)$ for every $\gamma\in\Gamma^{\dagger}_{1}$. Similarly, $\gamma(y)y^{-\chi_{2}(\gamma)}$ is contained in $\Pi(0,2)$ for every $\gamma\in\Gamma^{{\dagger}}_{2}$.
@@
We only prove the first assertion. Since $\Gamma^{{\dagger}}_{1}$ acts $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-linearly on $\Pi(1,0)/\Pi(2,0)$, which is a free $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-module generated by $x$, the action of $\gamma$ is given by a scalar multiplication by an element of $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]^{\times}$. However, since we have $0\neq z=[y,x]=(y-1)x$ in $\Pi(1,0)/\Pi(2,0)$ and $\gamma(z)=z^{\chi_{1}(\gamma)}$, such a scalar must coincide with $\chi_{1}(\gamma)$ as desired. ∎
@@
Let $\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}_{\geq 0}^{2}\setminus\{(0,0)\}$. Then the action of $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ on $\operatorname{Gr}^{\boldsymbol{m}}_{1}\Pi$ is given by the character $\chi_{1}^{m_{1}}$. Similarly, the action of $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}$ on $\operatorname{Gr}^{\boldsymbol{m}}_{2}\Pi$ is given by $\chi_{2}^{m_{2}}$.
@@
Fix an arbitrary element $\gamma\in\Gamma^{{\dagger}}_{1}$. We show that the action of $\gamma$ on $\operatorname{Gr}^{\boldsymbol{m}}_{1}\Pi$ is given by multiplication by $\chi_{1}^{m_{1}}(\gamma)$ by induction on $|\boldsymbol{m}|$. If $|\boldsymbol{m}|=1$, the assertion follows from (the proof of) Lemma 3.12. In general, the graded quotient $\operatorname{Gr}^{\boldsymbol{m}}_{1}\Pi$ is generated (as a $\mathbb{Z}_{p}[[\Pi/\Pi(1,0)]]$-module) by the image of commutator maps
$\cdot\;,\;\cdot\;]\colon\operatorname{Gr}^{\boldsymbol{m}^{\prime}}_{1}\Pi\times\operatorname{Gr}^{\boldsymbol{m}^{\prime\prime}}_{1}\Pi\to\operatorname{Gr}^{\boldsymbol{m}}_{1}\Pi,$
where $\boldsymbol{m}^{\prime}=(m_{1}^{\prime},m_{2}^{\prime})$ and $\boldsymbol{m}^{\prime\prime}=(m_{1}^{\prime\prime},m_{2}^{\prime\prime})$ satisfy $\boldsymbol{m}^{\prime}+\boldsymbol{m}^{\prime\prime}=\boldsymbol{m}$. Since this pairing is bilinear, for every $(\tau^{\prime},\tau^{\prime\prime})\in\operatorname{Gr}^{\boldsymbol{m}^{\prime}}_{1}\Pi\times\operatorname{Gr}^{\boldsymbol{m}^{\prime\prime}}_{1}\Pi$, we have
@@
$i_{\boldsymbol{m},1}\colon\operatorname{Gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}\to\operatorname{Gr}^{\boldsymbol{m}+(1,0)}_{1}\Pi(\chi_{1}^{-1})\oplus\operatorname{Gr}^{\boldsymbol{m}+(0,1)}_{1}\Pi$
is compatible with the action of $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$, where $\operatorname{Gr}^{\boldsymbol{m}+(1,0)}_{1}\Pi(\chi_{1}^{-1})$ denotes the $\chi_{1}^{-1}$-twist of $\operatorname{Gr}^{\boldsymbol{m}+(1,0)}_{1}\Pi$. Similarly, the homomorphism
$i_{\boldsymbol{m},2}\colon\operatorname{Gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}\to\operatorname{Gr}^{\boldsymbol{m}+(1,0)}_{2}\Pi\oplus\operatorname{Gr}^{\boldsymbol{m}+(0,1)}_{2}\Pi(\chi_{2}^{-1})$
is compatible with the action of $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}$. In particular, the action of $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ (resp. $\Gamma^{{\dagger}}_{2}/F^{(0,1)}\Gamma^{{\dagger}}$) on $\operatorname{Gr}^{\boldsymbol{m}}_{1}\tilde{\Gamma}$ (resp. $\operatorname{Gr}^{\boldsymbol{m}}_{2}\tilde{\Gamma}$) is given by $\chi_{1}^{m_{1}}$ (resp. $\chi_{2}^{m_{2}}$).
@@
Before proving the assertion, we note that the group $\Gamma^{{\dagger}}_{1}/F^{(1,0)}\Gamma^{{\dagger}}$ acts on the graded quotients $\operatorname{Gr}^{\boldsymbol{m}}_{1}\Pi$ by $\chi_{1}^{m_{1}}$ by Lemma 3.13. Take an arbitrary element $\gamma$ of $\Gamma^{{\dagger}}_{1}$. We prove the first assertion by computing the image of $\gamma\cdot f$ under the map $i_{\boldsymbol{m},1}$ for an arbitrary $f\in F^{\boldsymbol{m}}\tilde{\Gamma}$.
@@
Since $\alpha\coloneqq y^{-1}\gamma^{-1}(y)$ is contained in $\Pi(2)=\Pi(\boldsymbol{1})$, we have